OpenStudy (anonymous):

Given the polynomial 2x3 + 18x2 - 18x - 162, what is the value of the coefficient 'k' in the factored form? 2x3 + 18x2 - 18x - 162 = 2(x + k)(x - k)(x + 9)

OpenStudy (campbell_st):

so look at the constant term \[-162 = 2 \times k \times (-k) \times 9\] just solve for k... hope this helps

OpenStudy (anonymous):

I don't know how

OpenStudy (campbell_st):

ok... so you have \[-162 = -18k^2\] solve for k...

OpenStudy (campbell_st):

so divide both sides of the equation by -18 then take the square root

OpenStudy (anonymous):

3=k? @campbell_st

OpenStudy (campbell_st):

well its + or - 3... but it works for your problem as 3..

OpenStudy (anonymous):

Okay Thank You so much!!!

OpenStudy (anonymous):

@Monster1234 help

OpenStudy (anonymous):

hold on a sec while I work it out @Law&Order

OpenStudy (anonymous):

thx ;)

OpenStudy (anonymous):

i know this 2(x^3+9x^2-9x-81)

OpenStudy (anonymous):

do you want the answer to that?

OpenStudy (anonymous):

yeah i guess

OpenStudy (anonymous):

k